OpenStudy (anonymous):

The volume of a cone varies jointly with its height and the square of its base radius. A cone has a base radius of 4ft, height of 6 ft, and volume 100.48(cubed feet). What is volume of a cone with a height of 3ft, and a base radius of 3ft.

OpenStudy (lgbasallote):

since this is a proportion...i'll answer that way.. \[\huge V _{cube} = khr ^{2}\] <----k is the proportionality constant 100.48 = k(6)(4)^2 100.48 = k (6)(16) 100.48 = k (96) divide both sides... k = 1.046 so.... V = (1.046)(3)(3)^2 ^k ^h ^r i leave the rest to you yes?

OpenStudy (anonymous):

thanks a ton!

OpenStudy (lgbasallote):

<tips hat>